### v1

@@ V.3.3 Emulation in $(\mathrm{Re}\,E,\mathrm{Im}\,E,E_{rel},\lambda,\lambda_{4})$
In contrast to the previous emulations in Sec. V.3.1 and V.3.2, we need to vary the sources to emulate the particle-dimer elastic scattering amplitudes. Recall, for each on-shell scattering calculation or emulation, the $E_{rel}$ parameter in the sources is related to the $E$ variable via $E=E_{rel}-B_{2}$. Note $B_{2}$ is a function of $\lambda$. However, in training calculations, we only compute at complex $E$s. Therefore, the $E_{rel}$ and $E$ parameters are treated separately at the emulator off-line training stage. Meanwhile, the $\lambda$ and $\lambda_{4}$ parameters in the $H$ and sources are treated as the same.
@@
Therefore, we sample the training points in the $(\mathrm{Re}\,E,E_{rel},\lambda,\lambda_{4})$ space using LHS while fixing $\mathrm{Im}\,E$ to $3$ or $10$ MeV. We then perform emulations in the entire $(\mathrm{Re}\,E,\mathrm{Im}\,E,E_{rel},\lambda,\lambda_{4})$ space, including extrapolating training results at complex $E$s to the real energies.
@@
Figure 28: Emulation in the $(E,E_{rel},\lambda,\lambda_{4})$ space for the partilce-dimer scattering amplitude. We show the tests of the emulations at 16 cases with different $(B_{2},\lambda_{4})$ values. From left to right, the emulation interpolates for the $\lambda$ variable in the first two columns while it extrapolates in the last two columns. There are four rows, with each row having two sub-rows. From top to bottom, $\lambda_{4}$ values increase. The top sub rows plot elastic scattering observables, including the elastic scattering phase shift, normalized to $\pi$, and the modulus of the $S$ matrix offset by $-1$. The emulation results are compared to the exact results. The lower sub-row shows the relative errors for the full $T$-matrix, and its non-Born piece $T_{NB}$.
@@
Figure 29: Emulation in the $(E,E_{rel},\lambda,\lambda_{4})$ space for the particle-dimer scattering amplitude. We test the emulation performance in a grid in the ($\lambda,\lambda_{4}$) space with 250 points; for each ($\lambda,\lambda_{4}$), we pick an even mesh with 128 points for the $E_{rel}$ variable. The top two panels show the relative errors for emulating $T_{NB}$ vs $E_{rel}$ for different $B_{2}$ values but then averaged over $\lambda_{4}$. $E=E_{rel}-B_{2}$ is set there. Similar errors are plotted in the bottom panels but with $E=E_{rel}-B_{2}+10\,i$.
@@
In Fig. 28, we show the detailed emulation results for a sample of 16 test points with different combinations of $B_{2}$ and $\lambda_{4}$. There are four rows. Within each row, the panels in the top sub-row show rescaled scattering phase shifts ($\delta/\pi$) and the shifted elasticities ($\eta-1$) vs. particle-dimer scattering energy $E_{rel}$. The elastic scattering phase shift $\delta$ and the inelasticity $\eta$ are derived from the $S$ matrix via $\log(S)/2i=\delta+i\delta_{I}$ and $\eta\equiv e^{-2\delta_{I}}-1$. The emulation errors, including for the non-Born $T$-matrix and full $T$-matrix, are plotted in the bottom sub-row panels. We use two different meshes for the horizontal axis, one between 00 and $15$ and the other above $15$, to show the details around the particle-dimer threshold and, at the same time, present the trends up to 60 MeV.
@@
Again, the errors are generally tiny but grow toward the near-threshold region. At small $E_{rel}$, the $\eta$ emulation errors can reach $10^{-2}$ to $10^{-1}$, when $\mathrm{Im}\,E^{tr}_{\alpha}=10$ MeV. With $\mathrm{Im}\,E^{tr}=3$ MeV, those errors are suppressed significantly, as inferred in the sub rows. As a side note, in the last column with $B_{2}=1.5$ MeV (an extrapolation case), we see the signs of the virtual states can be seen from emulated phase shift results.
@@
In the sub rows, we notice enhancements of emulation errors for the $T$-matrix compared to the errors for the non-Born term due to strong cancellations between the Born and non-Born terms in their contribution to the full $T$-matrix. Across the table, the shallow dimer scattering errors are more significant than those for other dimers.
@@
To have a better understanding of the emulation performance, we systematically checked the two emulators by sampling a large number of test points. In the sampling, $\lambda$ and $\lambda_{4}$ values are distributed in an even grid in their two-dimensional space (250 grid points in total), and $E_{rel}$ is sampled independently from an even mesh (128 points).
@@
In the top row plots, we look at the emulations of the on-shell $T$-matrix, and thus $E$ is determined by $E_{rel}$ according to $E=E_{rel}-B_{2}$. The panels show the relative emulation errors for the emulated scattering amplitudes vs $E_{rel}$ for a series of $B_{2}$ values. Each curve corresponds to (with a fixed $B_{2}$) the mean of a result sample with different $\lambda_{4}$ values. We see dependence of the errors on $B_{2}$ value and $E_{rel}$, as we noticed in Fig. 28. The errors increase when approaching the particle-dimer threshold or reducing the $B_{2}$ value. Again, reducing $\mathrm{Im}\,E^{tr}_{\alpha}$ decreases the emulation errors in the region close to the particle-dimer threshold for all the $B_{2}$ values.
@@
In the bottom panels, we set $E$ according to $E=E_{rel}-B_{2}+10\,i$ so that we can probe the emulation performance for interpolating complex-$E$ $\mathcal{A}$ in the space of ($\lambda$, $\lambda_{4}$). Only two curves with the smallest and largest $B_{2}$ values are shown; the error curves fall between the two for the other $B_{2}$ values. Not surprisingly, these errors are systematically smaller than the extrapolation errors shown in the top panels. The errors increase slightly towards the particle-dimer threshold and higher energies outside the training-energy range. The further implication from the bottom panel is discussed in Sec. VI.
@@ VI Potential couplings with existing methods
The CEE developed here can analytically continue $\mathcal{A}$ from one region of the complex $E$ plane to the other. Therefore, when a particular Schrödinger equation solver or method is coupled with the CEE tool, i.e., applying the solver to obtain those training solutions, the reach of the solver is expanded in the complex $E$ plane.
@@
Specifically, for solvers based on Hermitian $H$ matrices (e.g., Harmonic oscillator basis methods), the new region they can access via CEE includes the real axis and the region below. I.e., the continuum physics can be computed using those solvers. As for the existing NHQM methods (e.g., the complex scaling method), they can solve the training equations with $E$ on the real axis or even below as long as $E$ is above their rotated BCTs. The CEE can help these methods reach the region to the left of these BCTs, i.e., rotate their BCTs even further away from the real axis.
@@
The CERPE tool can further expand the functionality of the existing solvers, as the emulator provides potential users easy access to these solvers to explore the parameter space of the solved model quickly. The applications have been mentioned in Sec. I and are thus not repeated here.
@@
It is worth mentioning that Ref. Liu et al. (2024) recently developed an emulator specifically for two-body coupled-channel scattering calculations based on the complex scaling method. However, the emulation in the complex $E$ plane and the spectrum emulation were not studied there.
@@
Another group of continuum physics methods, the CE and LIT methods, can also be coupled with our CERPE. Like us, these methods make use of the Schrödinger equation solutions at complex energies and extrapolate the results to real energy axis using their specific procedures. It is the complex energy calculation component that can be emulated using our CERPE method, which can significantly reduce the computing cost of exploring the parameter space of these calculations. See the discussion around the lower panels in Fig. 29 for the performance of the CERPE emulations at complex energies.
@@
The CE method has been primarily applied in the few-body studies Schlessinger and Schwartz (1966); Schlessinger (1968a, b); McDonald and Nuttall (1969); Uzu et al. (2003); Deltuva and Fonseca (2012, 2013a, 2013b, 2014). Although they have been focused on computing scattering amplitudes by solving integral equations, the emulations developed in this work can still apply because the wave functions or Faddeev components can be computed from these amplitudes and vice versa. Or, we can use our CERPE directly to emulate the integral equations, which are linear equations as well. Through either of these two routes, we can emulate the complex energy amplitudes in the space of the $\bm{\theta}$ parameters. These complex energy amplitudes can then be extrapolated to the real energy axis using the continuous fraction method, a rational approximation used extensively in the CE calculations.
@@
The LIT method Efros (1985a); Efros et al. (1994, 2007); Orlandini et al. (2014); Sobczyk et al. (2021, 2024); Bonaiti et al. (2024), which has been applied mostly for computing response functions [^21] Note Ref. Efros et al. (2007) pointed out ways to compute other amplitudes beyond response functions using the LIT type method, connect $R(E)\equiv-\mathrm{Im}\,\mathcal{A}$ at complex energy $E$ to $R(E_{0})$ at real energy $E_{0}$ via an integral relationship:
[…]
@@
The $R(E_{0})$ as a function of $E_{0}$ can then be extracted from $R(E)$ via solving a typical inverse problem Efros et al. (2007). Existing studies computed the $R(E)$ data using nuclear structure methods, such as the Hyperspherical harmonics basis and coupled cluster methods. The $R(E)$ calculations are the computationally expensive component. However, these calculations can be emulated in the input parameter space using CERPE. Then, by solving the inverse problem, we attain emulations of $R(E_{0})$ in $\bm{\theta}$. Moreover, the same strategy can emulate Green’s Function-based optical potential calculations Rotureau et al. (2017); Burrows et al. (2024).
@@ VII Summary
This work arrives at a remarkable conclusion concerning a parametrized inhomogeneous Schrödinger equation for a finite quantum system. We can efficiently emulate the continuum spectra and the associated $\mathcal{A}$ by adopting the RBM and treating the complex $E$ variable and other parameters in the equation in a whole parameter space. The spectra, as projected in a low-dimensional subspace and thus compressed, indicate that the method belongs to the NHQM category if the non-Hermitization is carried out by properly setting up the complex training energies. Therefore, this study adds a new NHQM method for studying continuum physics. Moreover, it offers a new understanding of related LIT and CE methods from the perspective of the NHQM framework. It also unifies them with other NHQM methods, such as the complex scaling method.
@@
We also see a close connection between the emulation in the complex $E$ plane (i.e., CEE) and the recent new understanding of the (near)-optimal rational approximation of univariate function with branch points, in terms of pole distributions (i.e., spectrum pattern) and the error scaling with the number of poles (i.e., the effective dimension of the emulator subspace, $N_{\mathrm{eff}}$).
@@
On the computational advantage, the CEE method allows us to extract continuum physics based on bound-state-like training calculations. For example, we could use the harmonic oscillator basis method to compute resonance without modifying the underlying $H$, which is needed in the other NHQM methods. The CERPE expands the CEE further by interpolating (and extrapolating) CEE in the space of other parameters in the inhomogeneous Schrödinger equation. The potential applications of the CERPE are immense, as discussed in Secs. I and VI.
@@
We also suggest a rudimentary test of a NHQM calculation, as discussed at the end of Sec. II.3, by inspecting its eigenenergy distribution (spectrum) in the complex energy plane. The physical and DC states must be well separated so that the observable calculation, such as $\mathcal{A}$, makes physical sense.
@@
However, the puzzle concerning the bias of the emulators developed here, as mentioned at the end of Sec. I, is still lingering. How does a small-matrix-based emulation work better than a large-Hermitian-matrix-based high-fidelity calculation for continuum physics, even though the emulators are trained by high-fidelity calculations? We go back to Fig. 1 and focus on the region above the positive real axis, i.e., $E_{0}+i\eta$ with $E_{0}$ being real and $\eta$ as a positive but small number. The difference between the finite $R_{IR}$ results and the $R_{IR}\to\infty$ limit shows at $R_{IR}\to\infty$, $\mathcal{A}$ changes in a much smoother way than the variations at finite $R_{IR}$. I.e., the finite $R_{IR}$ boundary condition changes $\mathcal{A}$ in a dramatic way  [^22] From the time-dependent picture, the dynamics of a wave packet scattering off a potential center and then into infinity is simpler than a wave packet scattering back and forth inside a finite spatial volume.. From the perspective of rational approximations, to describe that smooth transition, we need to move those BCT poles away from the real axis. Our discussion of (near)-optimal rational approximation in Sec. II.3 shows the DC states can be efficiently discretized while the approximation still reproduces $\mathcal{A}$, including in the near-threshold region. This explains why our emulators are well suited for describing the physical $\mathcal{A}$ instead of the $\mathcal{A}$ at finite $R_{IR}$.
@@
Undoubtedly, more studies need to be done in the future. The version of the RBM method, specifically the component relying on the variational approach for solving linear equations, could be further improved by employing better projection approaches or better test function space. The regularization, for which we use a truncation on the SVD singular values, could become more sophisticated by incorporating prior information. With these improvements, our emulators could tolerate significant numerical errors in the training calculations.
@@
The emulator developed here could potentially be transformed into a data-driven emulator by discovering the emulator from the training data $\mathcal{A}(E^{tr}_{\alpha},\bm{\theta}^{tr}_{\alpha})$ without projecting the underlying equation system. This would generalize the so-called parameterized matrix model (PMM) Cook et al. (2024), a recently proposed machine learning platform based on Hermitian or Unitary matrices, to a platform based on general matrices. Such a data-driven emulator would also generalize the (near)-optimal rational approximation to a multivariate approximation.

### v2

@@ V.3.3 Emulation in $(\mathrm{Re}\,E,\mathrm{Im}\,E,E_{\mathrm{rel}},\lambda,\lambda_{4})$
In contrast to the previous emulations in Sec. V.3.1 and V.3.2, I need to vary the sources to emulate the particle-dimer elastic scattering amplitudes. Recall that, for each on-shell scattering calculation or emulation, the $E_{\mathrm{rel}}$ parameter in the sources is related to the $E$ variable via $E=E_{\mathrm{rel}}-B_{2}$. Note $B_{2}$ is a function of $\lambda$. However, in training calculations, I only compute at complex $E$s. Therefore, the $E_{\mathrm{rel}}$ and $E$ parameters are treated separately at the emulator off-line training stage. Meanwhile, the $\lambda$ and $\lambda_{4}$ parameters in the $H$ and sources are treated as the same.
@@
In short, I sample the training points in the $(\mathrm{Re}\,E,E_{\mathrm{rel}},\lambda,\lambda_{4})$ space using LHS while fixing $\mathrm{Im}\,E$ to $3$ or $10$ MeV. Emulations are then performed in the entire $(\mathrm{Re}\,E,\mathrm{Im}\,E,E_{\mathrm{rel}},\lambda,\lambda_{4})$ space, including extrapolating training results at complex $E$s to the real energies.
@@
Figure 25: Emulation in the $(E,E_{\mathrm{rel}},\lambda,\lambda_{4})$ space for the partilce-dimer scattering amplitude. I show the tests of the emulations at 16 cases with different $(B_{2},\lambda_{4})$ values. From left to right, the emulation interpolates for the $\lambda$ variable in the first two columns while it extrapolates in the last two columns. There are four rows, with each row having two sub-rows. From top to bottom, $\lambda_{4}$ values increase. The top sub rows plot elastic scattering observables, including the elastic scattering phase shift, normalized to $\pi$, and the modulus of the $S$ matrix offset by 1-1. The emulation results are compared to the exact results. The lower sub-row shows the relative errors for the full $T$-matrix, and its non-Born piece $T_{\mathrm{nb}}$.
@@
Figure 26: Emulation in the $(E,E_{\mathrm{rel}},\lambda,\lambda_{4})$ space for the particle-dimer scattering amplitude. I test the emulation performance in a grid in the ($\lambda,\lambda_{4}$) space with 250 points; for each ($\lambda,\lambda_{4}$), an even mesh with 128 points is picked for the $E_{\mathrm{rel}}$ variable. The top two panels show the relative errors for emulating $T_{NB}$ vs $E_{\mathrm{rel}}$ for different $B_{2}$ values but then averaged over $\lambda_{4}$. $E=E_{\mathrm{rel}}-B_{2}$ is set there. Similar errors are plotted in the bottom panels but with $E=E_{\mathrm{rel}}-B_{2}+10\,i$.
@@
Figure 25 shows the detailed emulation results for a sample of 16 test points with different combinations of $B_{2}$ and $\lambda_{4}$. There are four rows. Within each row, the panels in the top sub-row show rescaled scattering phase shifts ($\delta/\pi$) and the shifted elasticities ($\eta-1$) vs. particle-dimer scattering energy $E_{\mathrm{rel}}$. The elastic scattering phase shift $\delta$ and the inelasticity $\eta$ are derived from the $S$ matrix via $\log(S)/2i=\delta+i\delta_{I}$ and $\eta\equiv e^{-2\delta_{I}}-1$. The emulation errors, including for the non-Born $T$-matrix and full $T$-matrix, are plotted in the bottom sub-row panels. I use two different meshes for the horizontal axis, one between 0 and $15$ and the other above $15$, to better illustrate the details around the particle-dimer threshold and, at the same time, present the trends up to 60 MeV.
@@
Again, the errors are generally tiny but grow toward the near-threshold region. At small $E_{\mathrm{rel}}$, the $\eta$ emulation errors can reach $10^{-2}$ to $10^{-1}$, when $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}=10$ MeV. With $\mathrm{Im}\,E^{\mathrm{tr}}=3$ MeV, those errors are suppressed significantly, as inferred in the sub rows. As a side note, in the last column with $B_{2}=1.5$ MeV (an extrapolation case), one can see the signs of the virtual states can be seen from emulated phase shift results.
@@
In the sub rows, I notice enhancements of emulation errors for the $T$-matrix compared to the errors for the non-Born term due to strong cancellations between the Born and non-Born terms in their contribution to the full $T$-matrix. Across the table, the shallow dimer scattering errors are more significant than those for other dimers.
@@
To have a better understanding of the emulation performance, I systematically checked the two emulators by sampling a large number of test points. In the sampling, $\lambda$ and $\lambda_{4}$ values are distributed in an even grid in their two-dimensional space ($250$ grid points in total), and $E_{\mathrm{rel}}$ is sampled independently from an even mesh ($128$ points).
@@
In the top row plots of Fig. 26, I look at the emulations of the on-shell $T$-matrix, and thus $E$ is determined by $E_{\mathrm{rel}}$ according to $E=E_{\mathrm{rel}}-B_{2}$. The panels show the relative emulation errors for the emulated scattering amplitudes vs $E_{\mathrm{rel}}$ for a series of $B_{2}$ values. Each curve corresponds to (with a fixed $B_{2}$) the mean of a result sample with different $\lambda_{4}$ values. One sees dependence of the errors on $B_{2}$ value and $E_{\mathrm{rel}}$, as was noticed in Fig. 25. The errors increase when approaching the particle-dimer threshold or reducing the $B_{2}$ value. Again, reducing $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}$ decreases the emulation errors in the region close to the particle-dimer threshold for all the $B_{2}$ values.
@@
In the bottom panels, the value of $E$ is set according to $E=E_{\mathrm{rel}}-B_{2}+10\,i$ so that I can probe the emulation performance for interpolating complex-$E$ $\mathcal{A}$ in the space of ($\lambda$, $\lambda_{4}$). Only two curves with the smallest and largest $B_{2}$ values are shown; the error curves fall between the two for the other $B_{2}$ values. As expected, these errors are systematically smaller than the extrapolation errors shown in the top panels. The errors increase slightly towards the particle-dimer threshold and higher energies outside the training-energy range. The implication from the bottom panel is further discussed in Sec. VI.
@@ VI Potential couplings with existing methods
The CEE developed here can analytically continue $\mathcal{A}$ from one region of the complex $E$ plane to the other. Therefore, when a particular Schrödinger equation solver or method is coupled with CEE—via applying the solver to obtain the training solutions required in CEE, the reach of the solver is expanded in the complex $E$ plane.
@@
Specifically, for solvers based on Hermitian $H$ matrices (e.g., Harmonic oscillator basis methods), the new region they can access via CEE includes the real axis and the region below. In this way, the continuum physics can be computed using such solvers. As for the existing NHQM methods (e.g., the complex scaling method), they can solve the training equations with $E$ on the real axis or even below it as long as $E$ is above their rotated BCTs. The CEE can help these NHQM methods reach the region to the left of these BCTs, effectively rotating their BCTs even further away from the real axis.
@@
The CERPE tool can further expand the functionality of all these solvers, by providing potential users easy access to the solvers to explore their predictions in the input-parameter spaces. The applications have been mentioned in Sec. I and are thus not repeated here.
@@
It is worth mentioning that Ref. [35] recently developed an emulator specifically for two-body coupled-channel scattering calculations based on the complex scaling method. However, the emulation in the complex $E$ plane and the spectrum emulation were not studied there.
@@
Another group of continuum physics methods, the CE and LIT methods, can also be coupled with the CERPE. Like the CEE, these methods make use of the Schrödinger equation solutions at complex energies and extrapolate the results to real energy axis using their specific procedures. It is the complex-energy calculation component that can be emulated using the CERPE method, significantly reducing the computing cost of this component when exploring the parameter space. The excellent performance of the CERPE emulations at complex energies has been demonstrated in Fig. 26. See the discussion about this point around the figure.
@@
The CE method has been primarily applied in the few-body studies [17, 58, 59, 60, 61, 62, 63, 64, 65]. Although they have been focused on computing scattering amplitudes by solving integral equations, the emulations developed in this work can still apply because the wave functions or Faddeev components can be computed from these amplitudes and vice versa. Or, one can use the CERPE directly to emulate the integral equations, as a type of linear equations. Through either of these two routes, one can emulate the complex energy amplitudes in $\bm{\theta}$. These complex energy amplitudes can then be extrapolated to the real energy axis using the continuous fraction method, a rational approximation used extensively in the CE calculations.
@@
The LIT method [18, 66, 67, 68, 69, 70, 71], which has been applied mostly for computing response functions [^17] Note Ref. [67] pointed out ways to compute other amplitudes beyond response functions using the LIT type method, and reviewed various amplitude calculations of exclusive electromagnetic-current-induced disintegration of nuclei., connect $\mathrm{Im}\,\mathcal{A}(E)$ at complex energy $E$ to the same function evaluated just above the real-$E$ axis (i.e., at $E_{0}+i0^{+}$ with $E_{0}$ being real):
$\frac{\mathrm{Im}\,\mathcal{A}(E)}{\mathrm{Im}\,E}$ $=\frac{1}{\pi}\int dE_{0}\frac{\mathrm{Im}\,\mathcal{A}(E_{0}+i0^{+})}{|E-E_{0}|^{2}}\ ,$ (62)
which is conventionally expressed as:
$\langle\Psi|\Psi\rangle$ $=\int dE_{0}\frac{R(E_{0})}{|E-E_{0}|^{2}}\,.$ (63)
To see the equivalence of these two expressions, one note that (1) $\langle\Psi|\Psi\rangle$ (with $|\Psi\rangle$ as the solution of Eq. (2) at complex $E$) is the same as $\mathrm{Im}\,\mathcal{A}(E)/\mathrm{Im}\,E$ with $|{S}\rangle=|\tilde{{S}}\rangle$ and (2) the response function $R(E_{0})=-\mathrm{Im}\,\mathcal{A}(E_{0}+i0^{+})/\pi$.
@@
The response function $R(E_{0})$ as a function of $E_{0}$ can then be extracted from $\langle\Psi|\Psi\rangle$ via solving a typical inverse problem [67]. Existing studies computed the $\langle\Psi|\Psi\rangle$ data using nuclear structure methods, such as the Hyperspherical harmonics basis and coupled cluster methods [67, 68, 69, 70, 71]. These expensive calculations can now be emulated using CERPE. By solving the inverse problem, one then obtains emulations of $R(E_{0})$ in $\bm{\theta}$. Moreover, the same strategy can be used to emulate Green’s Function-based optical potential calculations [78, 79].
@@ VII Summary
This work arrives at a remarkable conclusion concerning a parametrized inhomogeneous Schrödinger equation for a finite quantum system. That is, one can efficiently emulate the continuum spectra and the associated $\mathcal{A}$ by adopting the RBM and treating the complex $E$ variable and other parameters in the equation in a whole parameter space. The spectra, as projected in a low-dimensional subspace and thus compressed, indicate that the method belongs to the NHQM category if the non-Hermitization is carried out by properly setting up the complex training energies. Therefore, this study adds a new NHQM method for studying continuum physics. Moreover, it offers a new understanding of related LIT and CE methods from the perspective of the NHQM framework. It also unifies them with other NHQM methods, such as the complex scaling method.
@@
I also see a close connection between the emulation in the complex $E$ plane (i.e., CEE) and the recent new understanding of the (near-)optimal rational approximation of univariate function with branch points, in terms of pole distributions (i.e., spectrum pattern) and the error scaling with the number of poles (i.e., the effective dimension of the emulator subspace, $N_{\mathrm{eff}}$).
@@
On the computational advantage, the CEE method allows people to extract continuum physics based on bound-state-like training calculations. For example, one could, in principle, use the harmonic oscillator basis method  [^18] Whether or not harmonic oscillator basais is optimal for this purpose is a separate issue. to compute resonance without modifying the underlying $H$, which is needed in the other NHQM methods. The CERPE expands the CEE further by interpolating (and extrapolating) CEE in the space of other parameters in the inhomogeneous Schrödinger equation. The potential applications of the CERPE are immense, as discussed in Secs. I and VI.
@@
I also suggest a rudimentary test of a NHQM calculation, as discussed at the end of Sec. II.3, by inspecting its eigenenergy distribution (spectrum) in the complex energy plane. The physical and DC states must be well separated so that the observable calculation, such as $\mathcal{A}$, makes physical sense.
@@
However, the puzzle concerning the bias of the emulators developed here, as mentioned at the end of Sec. I, is still lingering. How does a small-matrix-based emulation work better than a large-Hermitian-matrix-based high-fidelity calculation for continuum physics, even though the emulators are trained by high-fidelity calculations? To answer this question, one can go back to Fig. 1 and focus on the region above the positive real axis, i.e., $E_{0}+i\eta$ with $E_{0}$ being real and $\eta$ as a positive but small number. The difference between the finite $R_{IR}$ results and the $R_{IR}\to\infty$ limit shows at $R_{IR}\to\infty$, $\mathcal{A}$ changes in a much smoother way than the variations at finite $R_{IR}$. That is, the finite $R_{IR}$ boundary condition changes $\mathcal{A}$ in a dramatic way  [^19] In the time-dependent picture, the dynamics of a wave packet scattering off a potential center and then into infinity is simpler than a wave packet scattering back and forth inside a finite spatial volume.. From the perspective of rational approximations, to describe that smooth transition, one needs to move those BCT poles away from the real axis. The discussion of (near-)optimal rational approximation in Sec. II.3 shows the DC states can be efficiently discretized while the approximation still reproduces $\mathcal{A}$, including in the near-threshold region. This explains why my emulators are well suited for describing the physical $\mathcal{A}$ instead of its approximation based on a high-dimensional Hermitian $H$ matrix, when $E$ is real and above kinematic thresholds.
@@
Undoubtedly, additional investigations are required. The version of the RBM method, specifically the component relying on the variational approach for solving linear equations, could be further improved by employing other projection approaches or test function space. The regularization, for which I use a truncation on the SVD singular values, could become more sophisticated if relevant prior information is incorporated [115]. Greedy algorithms can be implemented to reduce the computing costs of trainings and provide another avenue of regularization. Based on these improvements, CEE and CERPE could tolerate significant numerical errors in the training calculations.
One also needs to explore avenues to deal with non-affine parameters, such as the $E_{\mathrm{rel}}$ variable, to further minimize the computational costs of the emulations. Moreover, it is necessary to expand the framework developed here to treat systems with the presence of significant Coulomb effects in their continuum physics, likely by following the procedures outlined in the existing LIT studies [67].
@@
Lastly, these emulators could be transformed into a data-driven type by calibrating them directly to the training data $\mathcal{A}(E^{\mathrm{tr}}_{\alpha},\bm{\theta}^{\mathrm{tr}}_{\alpha})$ without explicitly projecting the underlying equation system. This would generalize the so-called parameterized matrix model (PMM) [116], a recently proposed machine learning platform based on Hermitian or Unitary matrices, to deal with problems with continuous spectra. Such data-driven emulators would also generalize the (near-)optimal rational approximations for a univariate function to a method for multivariate functions.